Displacement vector A points due east and has a magnitude of 2.98 km. Displacement vector B points due north and has a magnitude of 1.54 km. Displacement vector C boints due west and has a magnitude of 1 km. Displacement vector D points due south and has a magnitude of 1.4 km. Find the magnitude and direction (relative to due east) of the resultant vector A +B+C+D. magnitude direction km ocounterdlockwise from due east Additional Materials section 1 8
Answer #1

A + C = 2.98 - 1 = 1.98 km due east

And B+ D = 1.54-1.40 = 0.14km due north

Now A+B+C+D = √(1.98^2 + 0.14^2) = 1.9849 = 1.985km

And angle @ = Cos^-1(1.98/1.9849) = 4.05°

=) @ = 4.05°